OpenStudy (anonymous):

Write expression as a single logarithm: 1/3 (lnx+lny)-4lnz so... do I distribute the 1/3 first?

OpenStudy (freckles):

sounds fabulous to me

OpenStudy (freckles):

then use power rule for log

OpenStudy (freckles):

\[r \log_b(x)=\log_b(x^r)\]

OpenStudy (freckles):

then use product and quotient rule for log

OpenStudy (anonymous):

properties of logarithms: \[a \ln (b)=\ln (b ^{a})\] \[\ln (a)+\ln (b)=\ln (ab)\]

OpenStudy (anonymous):

And one more property: \[\ln (a)-\ln (b)=\ln (\frac{ a }{ b })\]

OpenStudy (anonymous):

1/3lnx+1/3lny= 1/3lnxy?

OpenStudy (freckles):

oh yeah you can do that :p

OpenStudy (freckles):

but you are still going to have to use power rule

OpenStudy (anonymous):

@drayder It would probably be easier to start within the parentheses